# SeNMo: A Self-Normalizing Deep Learning Model for Enhanced Multi-Omics Data Analysis in Oncology

### v1

@@ Abstract
Multi-omics research has enhanced our understanding of cancer heterogeneity and progression. Investigating molecular data through multi-omics approaches is crucial for unraveling the complex biological mechanisms underlying cancer, thereby enabling more effective diagnosis, treatment, and prevention strategies. However, predicting patient outcomes through the integration of all available multi-omics data is still an under-study research direction. Here, we present SeNMo (Self-normalizing Network for Multi-omics), a deep neural network that has been trained on multi-omics data across 33 cancer types. SeNMo is particularly efficient in handling multi-omics data characterized by high-width (many features) and low-length (fewer samples) attributes. We trained SeNMo for the task of overall survival of patients using pan-cancer multi-omics data involving 33 cancer sites from the Genomics Data Commons (GDC). The training multi-omics data includes gene expression, DNA methylation, miRNA expression, DNA mutations, protein expression modalities, and clinical data. We evaluated the model’s performance in predicting patient’s overall survival using the concordance index (C-Index). SeNMo performed consistently well in the training regime, reflected by the validation C-Index of $0.76$ on GDC’s public data. In the testing regime, SeNMo performed with a C-Index of $0.758$ on a held-out test set. The model showed an average accuracy of $99.8\%$ on the task of classifying the primary cancer type on the pan-cancer test cohort. SeNMo proved to be a mini-foundation model for multi-omics oncology data because it demonstrated robust performance, and adaptability not only across molecular data types but also on the classification task of predicting the primary cancer type of patients. SeNMo can be further scaled to any cancer site and molecular data type. We believe SeNMo and similar models are poised to transform the oncology landscape, offering hope for more effective, efficient, and patient-centric cancer care.
Keywords Cancer  $\cdot$ Oncology  $\cdot$ Multi-Omics  $\cdot$ Pan-Cancer  $\cdot$ Machine Learning  $\cdot$ Deep Learning  $\cdot$ Survival.
@@ 1 Introduction
Cancer data has multiple modalities, each offering distinct but complementary views of the disease [1]. Radiological images reveal structural and functional anomalies, histopathology slides provide cellular and tissue-level detail, clinical and Electronic Health Records (EHR) data encapsulate patient history and treatment outcomes, and molecular data such as genomics, transcriptomics, proteomics, and metabolomics uncover the underlying biological mechanisms driving cancer progression and response to therapy [2, 3, 4, 5]. Studying cancer through a multimodal perspective is critical for comprehensive understanding and effective treatment strategies [6]. Additionally, multimodal approaches facilitate personalized medicine, enhance our ability to predict disease-related outcomes, and advance our understanding and treatment of cancer [7].
@@
Multimodal and Multiomics Data. The growth of molecular data has greatly advanced cancer research [5]. The emergence of high-throughput sequencing technologies supported by the development of sophisticated bioinformatics tools and computational algorithms has ushered in an era of “omics" [8]. Multi-omics is a subset of multimodal data that specifically refers to the integrated analysis of various molecular data modalities including genomics, transcriptomics, proteomics, and metabolomics [9]. Multi-omics provides a comprehensive view of the biological processes and molecular mechanisms underlying cancer [10]. By combining these different layers of molecular data, multi-omics transcends the limitations of single-omic studies, which might only provide a partial view of the disease, by illustrating how various molecular components (DNA mutations, protein expression, and RNA expression, etc) interact within the complex biological network of cancer [11].
Pan-cancer Perspective. Researchers have studied individual cancers as well as pan-cancer data. Although studying individual cancers has shown significant benefits in understanding specific pathways and therapeutic responses, the pan-cancer approach offers a broader, more systemic view that can accelerate breakthroughs applicable across multiple types of cancer. Pan-cancer studies have enabled the identification of commonalities and differences across various cancer types, leading to insights that may not be as evident when focusing on a single cancer type [12]. Pan-cancer setting has led to the identification of universal cancer vulnerabilities, detailed the landscape of pathway alterations leading to the development of cross-cancer diagnostic tools and treatments, and characterized genetic mutations and molecular profiles across thousands of tumors, revealing shared oncogenic pathways and mutation patterns across different cancers revealing new features with potential clinical utility [13, 14, 15, 12]. Similarly, pan-cancer studies have identified key molecular signatures that can predict response to immunotherapy across different tumor types, demonstrating the clinical relevance of pan-cancer approach [16, 17].
@@
Existing Landscape of Pan-cancer Multi-omics Analysis. Traditionally, multimodal, multi-omics, and pan-cancer studies are conducted through a variety of techniques and methods that leverage advanced computational tools, bioinformatics, statistical methods, machine learning, and deep learning models to integrate and interpret complex oncology datasets. The data integration techniques in multi-omics are generally categorized into supervised and unsupervised methods, which can be further sub-categorized into (1) feature extraction (selection, extraction, and dimensionality reduction), (2) feature engineering (transformation, reducing dimensionality, normalizing and simplifying data, reducing noise, and alignment), (3) network-based methods (patient similarity networks, patient-drug networks, drug-drug networks, etc.), (4) clustering (grouping similar samples, stratification, feature selection, and grouping biological modules), (5) factorization (decompose or factorize features, multiple kernel learning, Bayesian consensus, similarity network fusion, NMF), and (6) Deep Learning (CNNs, MLPs, RNNs, Transformers, GNNs, etc.) [18, 9, 19]. Deep learning, a subset of machine learning characterized by the use of neural networks with many layers, has dramatically transformed the study of high-width (many features), low-length (fewer samples) molecular data [20, 21]. With its inherent capacity to model complex, non-linear relationships and to handle vast datasets, deep learning has proven adept at uncovering patterns that elude traditional analysis. There are elaborate reviews in the existing literature that analyze different pan-cancer, multimodal, multi-omics works [22, 23, 24, 25, 9, 26, 27, 28]. Leveraging multi-omics data for improved cancer diagnosis, prognosis, and treatment stratification presents a transformative opportunity. Recent literature underscores the shift towards computational integration of heterogeneous biological data, revealing critical insights into cancer’s multifaceted nature. Here we examine seminal works that encapsulate the paradigm shift in multi-omics cancer studies.
@@
A notable advancement is the utilization of Self-Normalizing Neural Networks (SNNs) for pan-cancer classification, which, as demonstrated by the study on CNV data from The Cancer Genome Atlas (TCGA) lung adenocarcinoma (LUAD), ovarian (OV), liver hepatocellular carcinoma (LIHC), and breast (BRCA) cancers, underscores the importance of feature selection in managing high-dimensional data for effective disease categorization [29]. The SNN model trained to perform pan-cancer classification yielded superior accuracy and macro F1 scores over traditional algorithms like Random Forest [29]. Complementing the pan-cancer approach, an integrative analysis combining histology-genomic data via multimodal deep learning offered a broad-spectrum understanding of cancer biology [30]. With an extensive dataset from TCGA encompassing 14 cancer types, a deep learning multimodal fusion (MMF) model outperformed attention-based multiple-instance learning (AMIL) model and self-normalizing network (SNN), showcasing the benefits of integrative analytics over singular data type analyses [30]. Further emphasizing multi-omics data integration, DeepProg, an ensemble framework combined deep learning and machine learning for prognosis prediction [31]. By processing RNA-Seq, miRNA sequencing, and DNA methylation for 32 cancer types from TCGA, DeepProg excelled in survival subtype prediction and risk stratification [31].
@@
Another study identified novel subgroups with similar molecular characteristics by combining different models of machine learning and deep learning [32]. By reducing dimensions of multi-omics features (mRNA, miRNA, DNA methylation, protein expression) and applying various classifiers, this approach identified subgroups across 33 tumor types. The authors argued that the number of samples should be commensurate with the number of dimensions for better prediction power of a learning model [32]. Another study used four types of -omics data (gene expression, miRNA expression, protein expression, and DNA methylation) for two datasets (TCGA-BLCA, TCGA-LGG) to predict Progression-Free Interval (PFI) and Overall Survival (OS) through Multiview Factorization AutoEncoder [33]. The identification of pan-cancer prognostic biomarkers through an integrated multi-omics data (DNA methylation, gene expression, somatic copy number alteration, and microRNA expression) across 13 cancers shows the performance of statistical and bioinformatic methods in survival-related gene discovery [34]. The predictive capability of multi-omics data is further evidenced in non-small cell lung cancer (NSCLC) survival prediction, where the combination of five modalities, microRNA, mRNA, DNA methylation, long non-coding RNA (lncRNA) and clinical data, showed superior C-indices compared to individual modalities [35].
@@
The advantage of multimodal data fusion for survival prediction is quantified across various cancer stages and types, with FUSED models exhibiting better average C-index compared to various machine learning and bioinformatics methods [36]. This approach combined clinical features with genomic, transcriptomic, and proteomic data in oncological prognostics across 33 cancer types [36]. Deep learning-based clustering method called MCluster-VAEs predicted subtype discovery using multi-omics data (mRNA, miRNA, DNA methylation, CNA) across 32 cancer types, outperforming traditional methods [37]. Decoupled contrastive learning model, DEDUCE, used multi-head attention decoupled contrastive learning approach for subtype clustering through multi-omics data consisting of gene expression, DNA methylation, and miRNA expression, across five cancer types (BRCA, GBM, SARC, LUAD, STAD) [38]. The authors used multi-head attention encoder network for cancer subtype discovery [38].
@@
Challenges and Opportunities of Pan-Cancer Multi-Omics Analysis. Although good for the task at hand, the above-mentioned methods often struggle to fully capture the complexity and heterogeneity of cancer due to their inherent limitations in handling and interpreting vast, multidimensional datasets. Dimensionality reduction methods such as PCA or t-SNE can inadvertently discard subtle yet crucial biological nuances that might be pivotal for understanding disease mechanisms [39]. Learning-based dimensionality reduction methods, such as using deep learning models, lack the discriminating and interpreting ability of extracted features, lack consensus in the balance between the number of deep network layers and the number of layer neurons, and cannot handle or recover the missing data [39].
@@
Similarly, feature selection and learning-based feature engineering, despite their effectiveness in identifying key predictors within datasets, can introduce biases and result in models that are overfitted to specific features of the data used for training [40, 41]. This compromises such methods’ ability to perform well across different datasets or in real-world clinical settings [40, 41]. Furthermore, these methods frequently face challenges in terms of generalizability, as they may not perform consistently across diverse patient populations or varying biological conditions, limiting their utility in broader clinical practice. Thus, while these techniques are instrumental in advancing cancer research, their limitations highlight the need for more robust and generalizable framework that can more accurately predict end-points across different cancer types and data modalities.
@@
Recently, a new class of deep learning models called foundation models that comprise LLMs (Large Language Models) and VLMs (Vision-Language Models) have been trained on large multimodal data [42]. These models have demonstrated a strong ability to generalize well across different tasks when provided with ample and diverse training data [42]. Due to their extensive and varied datasets, these models capture a broad range of patterns and nuances, enabling them to apply learned knowledge flexibly and effectively across different contexts. The key conclusions from the success story of foundation models relevant to this study are:
@@
    Extensive Training Data: Foundation models are trained on massive datasets that encompass a wide spectrum of information across different domains and modalities. This extensive training helps the models develop a robust understanding of complex patterns and relationships within the data. For instance, models like GPT (from OpenAI) [43] and BERT (developed by Google) [44] have been shown to perform exceptionally well on a variety of natural language processing tasks, from translation to sentiment analysis, precisely because they have been exposed to large amounts of diverse textual data during training [45, 42].
@@
    Cross-Modal Learning: VLMs integrate information from both visual and textual sources, allowing them to develop a more comprehensive understanding of the world. For example, models like CLIP (from OpenAI) [46] and ViLBERT (Vision-and-Language BERT) [47] learn to correlate images with text, enabling them to perform tasks such as image captioning or visual question answering with high accuracy. This ability to process and synthesize information across different modalities enhances their flexibility and adaptability to new tasks that may not strictly resemble those they were originally trained on [45].
@@
    Generalization Across Tasks: The capability of these models to generalize is further evidenced by their performance across a range of tasks with minimal task-specific tuning [42]. For example, once trained, these models can often switch between tasks such as text classification, summarization, and even complex reasoning without needing extensive retraining. This adaptability is largely due to their training datasets’ comprehensive and diverse nature, which provides a rich background against which the models can evaluate new problems [42, 9, 45].
@@
The establishment of large-scale biological databases and data repositories, such as National Cancer Institute’s (NCI) The Cancer Genome Atlas (TCGA) [48] and Clinical Proteomic Tumor Analysis Consortium (CPTAC) [49], hold vast amounts of cancer data, especially multi-omics data, that can be readily used for disease analysis. Despite many efforts in building a foundation model for the omics data, the existing literature has no foundation model that has been trained on multi-omics pan-cancer data. scGPT is the foundation model trained for the single-cell sequencing data of 33 million cells [50]. SAMMS model has been trained on two cancer types (TCGA’s LGG and KIRC) using clinical (age, gender), gene expression, CNV, and miRNA, and WSI data [51]. RNA Foundation Model (RNA-FM) was trained on 23 million non-coding RNA sequences [52]. PATH-GPTOMIC used copy number variations (CNV), genomic mutations, bulk RNA Seq, and WSI data to predict survival outcomes for two datasets (TCGA-GBMLGG, TCGA-KIRC) [53]. The absence of a pan-cancer, multi-omics foundation model can be attributed to the complexity and heterogeneity of such data, lack of comprehensive datasets, specificity of current analytical methods, and large computational and resource constraints. To address these challenges, we consider a multi-omics, pan-cancer framework that involves only essential pre-processing steps. We propose a mini-foundation model called SeNMo (Self-Normalizing Deep Learning Model for Multi-Omics). SeNMo has been trained on six data modalities including clinical, gene expression, miRNA expression, DNA Methylation, DNA Mutations, and reverse-phase protein array (RPPA) expression data across 33 cancer types. We call it a mini-foundation model because the latent representations generated by SeNMo have not been evaluated for properties such as generalization, emergence, expressivity, scalability, and compositionally, which are essential traits for a model to be named as “foundation model" [54, 42]. We evaluated SeNMo’s generalization capability to tasks such as OS prediction and primary cancer classification. The rest of the evaluation is beyond the scope of this work and shall be pursued in future endeavors.
Predicting overall survival and accurately classifying cancer types are pivotal endpoints in cancer research and patient care. For patients, these predictions can inform treatment options, influence monitoring strategies, and guide clinical decision-making, thus impacting quality of life and survival prospects. For healthcare systems and researchers, the ability to predict outcomes enhances understanding of the disease, improves the design of clinical trials, and drives the development of new therapeutics. In essence, enhancing the accuracy of survival predictions and cancer classification through advanced computational methods not only stands to revolutionize the clinical approach to oncology but also embodies the patient-centered ethos that is central to modern medicine.
@@
Figure 1: Overview of the SeNMo framework. Data from public and private sources are collected and curated through the MINDS system [28]. The resulting cohorts of multimodal data is preprocessed and fed to the self-normalizing deep learning encoder network that learns the cross-modality, pan-cancer features. The learned encoder weights can be used for different downstream tasks such as predicting overall survival (OS), progression-free survival (PFS), or cancer classification using different prediction heads.

### v2

# Self-Normalizing Foundation Model for Enhanced Multi-Omics Data Analysis in Oncology

@@ Abstract
Multi-omics research has enhanced our understanding of cancer heterogeneity and progression. Investigating molecular data through multi-omics approaches is crucial for unraveling the complex biological mechanisms underlying cancer, thereby enabling more effective diagnosis, treatment, and prevention strategies. However, predicting patient outcomes through the integration of all available multi-omics data is still an under-study research direction. Here, we present SeNMo (Self-normalizing Network for Multi-omics), a foundation model that has been trained on multi-omics data across 33 cancer types. SeNMo is particularly efficient in handling multi-omics data characterized by high-width (many features) and low-length (fewer samples) attributes. We trained SeNMo for the task of overall survival of patients using pan-cancer multi-omics data involving 33 cancer sites from the Genomics Data Commons (GDC). The training multi-omics data includes gene expression, DNA methylation, miRNA expression, DNA mutations, protein expression modalities, and clinical data. SeNMo was validated on two independent cohorts: Moffitt Cancer Center and CPTAC lung squamous cell carcinoma. We evaluated the model’s performance in predicting patient’s overall survival using the concordance index (C-Index). SeNMo performed consistently well in the training regime, reflected by the validation C-Index of $0.76$ on GDC’s public data. In the testing regime, SeNMo performed with a C-Index of $0.758$ on a held-out test set. The model showed an average accuracy of $99.8\%$ on the task of classifying the primary cancer type on the pan-cancer test cohort. SeNMo demonstrated robust performance on the classification task of predicting the primary cancer type of patients. SeNMo further demonstrated significant performance in predicting tertiary lymph structures from multi-omics data, showing generalizability across cancer types, molecular data types, and clinical endpoints. We believe SeNMo and similar models are poised to transform the oncology landscape, offering hope for more effective, efficient, and patient-centric cancer care.
Keywords Cancer  $\cdot$ Oncology  $\cdot$ Multi-Omics  $\cdot$ Multimodal  $\cdot$ Pan-Cancer  $\cdot$ Machine Learning  $\cdot$ Foundation Model  $\cdot$ Survival.
@@ 1 Introduction
Across the cancer care continuum, from screening, diagnosis, treatment, to survivorship, vast amounts of standard-of-care data are collected from patients. In cancer research, the volume and diversity of data further expand, providing distinct and complementary views of the disease [1]. For instance, radiological images capture structural and functional information at the organ and sub-organ levels, histopathology slides offer morphological, cellular, and tissue-level insights, clinical and Electronic Health Records (EHR) encapsulate patient history, treatment plans, and outcomes, while molecular data—such as genomics, transcriptomics, proteomics, and metabolomics—reveal the underlying biological mechanisms driving cancer progression and treatment response [2, 3, 4, 5]. Studying cancer from a multimodal perspective is essential for comprehensive understanding and for developing effective, personalized treatment strategies [6, 7].
@@
Multimodal and multi-omics data. The advancement of technologies to record, process, and store molecular data has significantly propelled cancer research [5]. High-throughput sequencing technologies, along with sophisticated bioinformatics tools and computational algorithms, have ushered in an era of “omics" [8]. Multi-omics, a subset of multimodal data, specifically refers to the integrated analysis of various molecular modalities, including genomics, transcriptomics, proteomics, and metabolomics [9]. Multi-omics provides a comprehensive understanding of the biological processes and molecular mechanisms underlying cancer [10]. By combining different layers of molecular data, multi-omics transcends the limitations of single-omic studies, which often provide only a partial view of the disease. It illustrates how various molecular components, such as DNA mutations, protein expression, and RNA expression, interact within the complex biological network of cancer [11].
Pan-cancer perspective. Cancer research can be approached from two primary perspectives: individual cancer studies and pan-cancer studies. Individual cancer studies focus on a specific type of cancer, delving deep into its unique molecular and genetic characteristics, allowing for the development of highly targeted therapies and personalized treatment plans. Studying individual cancers has shown significant benefits in understanding specific pathways and therapeutic responses. Conversely, pan-cancer studies analyze commonalities and differences across multiple cancer types, uncovering shared molecular mechanisms and genetic alterations. This approach reveals broader patterns and potentially identifies universal biomarkers or therapeutic targets applicable across different cancers, enhancing our holistic understanding of the disease [12]. The pan-cancer perspective has uncovered universal cancer vulnerabilities, detailed pathway alterations for cross-cancer diagnostics and treatments, and revealed shared oncogenic pathways and mutation patterns, leading to new clinically useful insights [13, 14, 15, 12]. Furthermore, pan-cancer studies have identified key molecular signatures that can predict response to immunotherapy across diverse tumor types, demonstrating the wide-reaching clinical significance of the pan-cancer approach [16, 17]. In this article, we focus on the pan-cancer perspective, emphasizing its potential to generate overarching insights that could lead to more comprehensive and versatile cancer treatment strategies.
@@
Existing landscape of pan-cancer multi-omics analysis. Traditionally, multimodal, multi-omics, and pan-cancer studies have been conducted through a variety of techniques and methods that leverage advanced computational, bioinformatics, statistical, machine learning, and deep learning approaches to integrate and interpret complex oncology datasets. Data integration techniques in multi-omics are generally categorized into supervised, weakly supervised, and unsupervised methods. These methods can be further sub-categorized into (1) feature extraction (selection, extraction, and dimensionality reduction), (2) feature engineering (transformation, dimensionality reduction, data normalization, simplification, noise reduction, and alignment), (3) network-based methods (e.g., patient similarity networks, patient-drug networks, drug-drug networks), (4) clustering (e.g., grouping similar samples, stratification, feature selection, biological module grouping), (5) factorization (e.g., feature decomposition, multiple kernel learning, Bayesian consensus, similarity network fusion, non-negative matrix factorization), and (6) deep learning techniques (e.g., Convolutional Neural Networks (CNNs), Multilayer Perceptions (MLPs), Recurrent Neural Networks (RNNs), Transformers, Graph Neural Networks (GNNs)) [18, 9, 19, 20]. Deep learning, a subset of machine learning characterized by neural networks with many layers, has transformed the study of high-dimensional, low-sample molecular data [21, 22]. With its capacity to model complex, non-linear relationships and handle vast datasets, deep learning has proven adept at uncovering patterns that traditional statistical and machine learning models may not identify. Numerous reviews in existing literature provide in-depth analysis of various pan-cancer, multimodal, and multi-omics research efforts [23, 24, 25, 26, 9, 27, 28, 29].
@@
A significant advancement in the field is the use of self-normalizing neural networks for pan-cancer classification. A study leveraging copy number variation data from The Cancer Genome Atlas (TCGA) for lung adenocarcinoma (LUAD), ovarian cancer (OV), liver hepatocellular carcinoma (LIHC), and breast cancer (BRCA) demonstrated that feature selection is crucial for managing high-dimensional data in disease categorization [30]. The self-normalizing model for pan-cancer classification yielded superior accuracy and macro F1 scores compared to a traditional random forest algorithm [30]. Complementing this approach, an integrative analysis that combined histology-genomic data using multimodal deep learning provided broad-spectrum insights into cancer biology [31]. Using an extensive dataset from TCGA encompassing 14 cancer types, a deep learning multimodal fusion model outperformed an attention-based multiple-instance learning model and a self-normalizing network, demonstrating the benefits of integrative analytics over single data type analyses [31]. Emphasizing multi-omics data integration, DeepProg—an ensemble framework that combines deep learning and machine learning—achieved high performance in prognosis prediction [32]. By processing RNA-Seq, miRNA sequencing, and DNA methylation data for 32 cancer types from TCGA, DeepProg excelled in predicting survival subtypes and risk stratification [32].
@@
Khadirnaikar et al. identified novel subgroups with similar molecular characteristics by combining different machine learning and deep learning models [33]. By reducing the dimensionality of multi-omics features (e.g., mRNA, miRNA, DNA methylation, protein expression) and applying multiple classifiers, this approach successfully identified subgroups across 33 tumor types. The authors argued that the number of samples should be proportional to the number of features for optimal predictive power of a learning model [33]. Another study used four types of -omics data (gene expression, miRNA expression, protein expression, and DNA methylation) for two datasets (TCGA-BLCA, TCGA-LGG) to predict progression-free interval and overall survival (OS) through a multiview factorization autoencoder [34]. The identification of pan-cancer prognostic biomarkers using integrated multi-omics data (including DNA methylation, gene expression, somatic copy number alteration, and miRNA expression) across 13 cancers highlighted the power of statistical and bioinformatics methods for discovering survival-related genes [35]. The predictive capability of multi-omics data was also evident in non-small cell lung cancer survival prediction, where combining five modalities—miRNA, mRNA, DNA methylation, long non-coding RNA, and clinical data—resulted in a superior concordance index (C-Index) compared to individual modalities [36].
@@
The advantage of multimodal data fusion for predicting OS was quantified across various cancer stages and types, with fused models exhibiting higher average C-Index compared to machine learning and bioinformatics methods [37]. This approach combined clinical features with genomic, transcriptomic, and proteomic data in oncological prognostics across 33 cancer types [37]. A deep learning-based clustering method called MCluster-VAEs achieved superior performance in subtype discovery using multi-omics data (e.g., mRNA, miRNA, DNA methylation, CNA) across 32 cancer types [38]. The decoupled contrastive learning model DEDUCE employed a multi-head attention decoupled contrastive learning approach for subtype clustering through multi-omics data consisting of gene expression, DNA methylation, and miRNA expression across five cancer types (BRCA, GBM, SARC, LUAD, STAD) [39]. The authors of DEDUCE utilized a multi-head attention encoder network for cancer subtype discovery [39].
@@
Limitations of the State-of-the-art Methods. Although valuable for their intended tasks, the above-mentioned methods often struggle to fully capture the complexity and heterogeneity of cancer due to inherent limitations in handling and interpreting vast, multidimensional datasets. Dimensionality reduction methods such as principal component analysis or t-distributed stochastic neighbor embedding can inadvertently discard subtle yet crucial biological nuances that are pivotal for understanding disease mechanisms [40]. Learning-based dimensionality reduction methods, such as those utilizing deep learning, face challenges including limited discriminative and interpretive capabilities of extracted features, lack of consensus on the balance between the number of network layers and the number of neurons per layer, and limitations in handling or recovering missing data [40].
@@
Similarly, feature selection and learning-based feature engineering, despite being effective in identifying key predictors, can introduce biases and create models that are overly tailored to specific features within training datasets [41, 42]. This bias undermines generalizability across diverse datasets or real-world clinical settings [41, 42]. Additionally, these methods frequently face challenges in ensuring consistent performance across varied patient populations and biological conditions, limiting their broader clinical utility. Thus, while these techniques are instrumental in advancing cancer research, they underscore the need for more robust and generalizable frameworks capable of accurately predicting endpoints across diverse cancer types and data modalities.
@@
Recently, a new class of deep learning models called foundation models, which include large language models and vision-language models, have been introduced by training on large multimodal datasets [43, 44]. These models have demonstrated a strong ability to generalize across different tasks when provided with diverse and substantial training data [43]. By leveraging extensive and varied datasets, these models are able to capture a broad spectrum of patterns and nuances, allowing for flexible and effective application across different contexts. Key conclusions from the successes of foundation models relevant to this study are:
@@
    Extensive training data: Foundation models are trained on massive datasets encompassing diverse domains and modalities. This extensive training helps models develop a robust understanding of complex patterns and relationships within data. For example, Generative Pre-trained Transformers (GPT) [45] and Bidirectional Encoder Representations from Transformers (BERT) [46] have been shown to excel across various natural language processing tasks, from translation to sentiment analysis, due to exposure to large and varied textual datasets during training [44, 43].
@@
    Cross-modal learning: Vision-language models integrate visual and textual information, enabling comprehensive understanding. Models like Contrastive Language-Image Pre-Training (CLIP) [47] and Vision-and-Language BERT (ViLBERT) [48] correlate images and text, allowing them to perform tasks such as image captioning or visual question answering with high accuracy. This cross-modal processing and synthesizing capability enhances adaptability to new tasks beyond their original training scope [44].
@@
    Generalization across tasks: Foundation models exhibit impressive generalization across diverse tasks with minimal task-specific tuning [43]. Once trained, they can switch between tasks like text classification, summarization, and complex reasoning without extensive retraining. This adaptability is largely due to their training datasets’ comprehensive and diverse nature, which provides a rich background against which the models can evaluate new problems [43, 9, 44].
@@
The establishment of large-scale biological databases and data repositories, such as the National Cancer Institute’s TCGA [49] and the Clinical Proteomic Tumor Analysis Consortium (CPTAC) [50], hold vast amounts of multi-omics cancer data that are readily available for disease analysis. Despite numerous efforts, existing literature lacks a foundation model trained on multi-omics pan-cancer data. scGPT is a foundation model trained for single-cell sequencing data comprising 33 million cells [51]. The SAMMS model was trained on two cancer types (TCGA’s LGG and KIRC) using patient-level data (age, gender), gene expression, CNV, miRNA, and WSI [52]. The RNA Foundation Model (RNA-FM) was trained on 23 million non-coding RNA sequences [53]. PATH-GPTOMIC utilized CNV, genomic mutations, bulk RNA Seq, and WSI data to predict survival outcomes for two datasets (TCGA-GBMLGG, TCGA-KIRC) [54]. The absence of a pan-cancer, multi-omics foundation model can be attributed to challenges such as data complexity, heterogeneity, limited comprehensive datasets, specificity of analytical methods, and large computational demands. To address these challenges, we propose a multi-omics, pan-cancer framework with minimal preprocessing, introducing a foundation model called the ’Self-Normalizing Deep Learning Model for Multi-Omics’ (SeNMo). SeNMo has been trained on six data modalities, including clinical, gene expression, miRNA expression, DNA methylation, DNA mutations, and reverse-phase protein array (RPPA) expression data across 33 cancer types. We have evaluated SeNMo for generalization, scalability, emergence, expressivity, and compositionality, which are essential traits for a true foundation model [55, 43]. We evaluated SeNMo’s generalization capability to unseen datasets and across different tasks such as OS prediction, primary cancer classification, and tertiary lymph structures (TLS) ratio prediction. Figure 1 presents the overview of our framework.
This work offers the following contributions:
  1. 1.
    We present an oncology data analysis using molecular correlates of patient prognosis across 33 cancer types, addressing both disease-wide and individual patient levels.
  2. 2.
    We created a multi-omics, pan-cancer framework with minimal and essential preprocessing steps, eliminating the need for complex, custom-engineered methods, thereby allowing a greater focus on the learning aspect.
  3. 3.
    We developed a foundation model capable of generalizing across different tasks and to unseen data through fine-tuning.
  4. 4.
    Our findings indicate that MLP-based networks are highly susceptible to catastrophic forgetting. We demonstrate that fine-tuning should involve a fraction of the epochs ($\leq 30$), while adjusting the learning rate, weight decay, and dropout to fractionally update all layers of the trained model.
  5. 5.
    The SeNMo framework represents the first initiative to analyze 33 cancer types using six molecular data modalities: clinical data, gene expression, miRNA expression, DNA methylation, DNA mutations, and protein expression.
  6. 6.
    We present the first effort to predict tertiary lymph structures (TLS) ratio from multi-omic data only.
  7. 7.
    We provide the latent feature vectors (embeddings) learned by SeNMo as an open-access vector database system, HoneyBee, available through Hugging Face and GitHub.
@@
Figure 1: Overview of the SeNMo model. The data from public sources are collected using MINDS [29] and curated to develop the multimodal dataset for SeNMo training. MINDS is a metadata framework for fusing publicly available data sources like TCGA-GDC and UCSC Xena Portal into machine learning-ready format [49, 56, 29]. The dataset is preprocessed and fed to the self-normalizing deep learning encoder network that learns underlying sub-visual patterns from cross-modality, pan-cancer data. The learned encoder weights are later used for different downstream tasks (with or without fine-tuning), such as predicting the overall survival (OS), progression-free survival (PFS), cancer subtype classification, grading, or tertiary lymph structures (TLS) ratio.